# Spatial Proportional Hazards Model with Differential Regularization

### v1

@@ 5 Model Version for Areal data
6 Simulation Results
To demonstrate the performance of the proposed model, we present simulations on a horseshoe domain, as defined in Ramsay (2002); Wood et al. (2008). We use the spatial field $f(x,y)$, depicted in Figure 2(a) and defined as follows:
where $c$ is a constant chosen so that the minimum value of $f$ over the horseshoe domain is zero, approximately $c\simeq 1.77$, $r=0.5$ and $r_{0}=0.1$.
We randomly sample $n=100,200,300$ points, denoted by $\{p_{i}\}_{i=1}^{n}$, from the horseshoe domain, as illustrated in Figure 2(b). Figure 2(c) shows the corresponding Delaunay triangulation. Each sampled point is associated with an observation $x_{i}$ of a random variable $X_{i}$, which follows a standard normal distribution. Additionally, we observe $(y_{i},\delta_{i})$, where $y_{i}=\min(T_{i},C_{i})$ and $\delta_{i}=I(T_{i}\leq C_{i})$. The variable $T_{i}$ given $X_{i}=x$ follows a proportional hazards model with baseline hazard $\lambda_{0}(t)=1$ and $\beta_{0}=0.2$, as defined in equation (1), with the form $\lambda(t|x)=\lambda_{0}(t)\exp(x\beta_{0}+f(p_{i}))$. To introduce censoring, we set $C_{i}=1$ for each $i$, resulting in approximately 20% censored observations. We set $\mu=0.5,1,5,10$ in order to understand how the value of $\mu$ effects the estimation of the spatial field $f$ and of the regression coefficients vector of $\beta$. The finite elements method is applied using a basis of order $1$ constructed on the Toggle the table of contents Delaunay triangulation over the sampled locations.
(a) Spatial field $f(x,y)$ over the horseshoe domain
(b) Sampled points over the horseshoe domain
(c) Mesh based on the sampled points
We conduct $N=100$ simulations for each combination of $n$ and $\mu$. And example of the estimated spatial field in Figure 3. This figure compares the true values (Figure 3(a)) with the estimated values (Figure 3(b)), and reports the pointwise error. In table
(a) True Values
(b) Estimated Values
(c) Error Values
Figure 3: Comparison of true values, estimated values, and error values.
$n$ $\mu$ MSE of $f$ (sd) Bias of $\beta$ (sd)
100 0.5 0.168 (0.112) -0.008 (0.130)
1 0.168 (0.118) -0.007 (0.130)
3 0.177 (0.140) -0.006 (0.130)
5 0.186 (0.165) -0.006 (0.130)
10 0.190 (0.188) -0.007 (0.129)
200 0.5 0.157 (0.119) 0.008 (0.096)
1 0.160 (0.129) 0.007 (0.096)
3 0.168 (0.150) 0.007 (0.095)
5 0.180 (0.169) 0.007 (0.094)
10 0.185 (0.179) 0.008 (0.093)
300 0.5 0.141 (0.105) -0.003 (0.077)
1 0.142 (0.115) -0.002 (0.077)
3 0.143 (0.129) 0.000 (0.077)
5 0.144 (0.136) 0.000 (0.077)
10 0.148 (0.149) 0.000 (0.077)
Table 1: Simulation results for varying sample sizes ($n$) and smoothing parameters ($\mu$), showing the mean squared error (MSE) of the spatial field estimator $f$ and the bias of the parameter $\beta$. Standard deviations (sd) are reported in parentheses for both MSE and bias.
@@ 7 Empirical Application
//TODO
@@ 8 Conclusions
//TODO
@@ Appendix A Functional PIRLS Algorithm details
The maximizer $(\hat{\beta},\hat{f})$ of (4) must satisfy the following system of first-order equations:
$\begin{cases}\frac{\partial\mathcal{L}(\hat{\beta},\hat{f})}{\partial\beta_{k}}=0,\quad\forall k=1,\dots,q,\\ \lim_{\tau\to 0}\frac{1}{\tau}\left[\mathcal{L}(\hat{\beta},\hat{f}+\tau u)-\mathcal{L}(\hat{\beta},\hat{f})\right]-\mu m(u,\hat{f})=0,\quad\forall u\in\mathcal{F}.\end{cases}$ (12)
This system involves derivatives with respect to both parameters $\beta$ and $f$, where the derivative with respect to $f$ is the Gâteaux derivative in the direction of $u\in\mathcal{F}$.
For each $k=1,\dots,q$, the term $\frac{\partial\mathcal{L}(\hat{\beta},\hat{f})}{\partial\beta_{k}}$ is given by:
$\frac{\partial\mathcal{L}(\hat{\beta},\hat{f})}{\partial\beta_{k}}=\sum_{i:\delta_{i}=1}\left(x_{ik}-\frac{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)x_{jk}}{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)}\right).$ (13)
@@
$\mathcal{L}(\hat{\beta},\hat{f}+\tau u)-\mathcal{L}(\hat{\beta},\hat{f})$ $=\sum_{i:\delta_{i}=1}\left(tu(p_{i})-\log\frac{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{i})+tu(p_{i})\right)}{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)}\right)$
$=\sum_{i:\delta_{i}=1}\left(tu(p_{i})-\log\frac{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)(1+tu(p_{i})+o(t))}{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)}\right)$
$=\sum_{i:\delta_{i}=1}\left(tu(p_{i})-\log\left\{1+t\frac{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)u(p_{i})}{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)}+o(t)\right\}\right)$
$=\sum_{i:\delta_{i}=1}\left(tu(p_{i})-t\frac{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)u(p_{i})}{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)}+o(t)\right).$
@@
$\lim_{\tau\to 0}\frac{1}{\tau}\left[\mathcal{L}(\hat{\beta},\hat{f}+\tau u)-\mathcal{L}(\hat{\beta},\hat{f})\right]$ $=\sum_{i:\delta_{i}=1}\left(u(p_{i})-\frac{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)u(p_{i})}{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)}\right).$ (14)
@@
$\theta_{i}(\beta,f)$ $=\exp(x_{i}^{\top}\beta+f(p_{i})),$
@@
$s_{i}(\beta,f)$ $=\sum_{j=1}^{n}r_{ji}\theta_{j}(\beta,f).$
@@
$\frac{\partial\mathcal{L}(\beta,f)}{\partial\beta_{k}}$ $=\sum_{i=1}^{n}\delta_{i}x_{ik}-\sum_{i=1}^{n}\frac{\delta_{i}}{s_{i}(\beta,f)}\sum_{j=1}^{n}r_{ji}\theta_{j}(\beta,f)x_{jk}$
$=\sum_{i=1}^{n}\delta_{i}x_{ik}-\sum_{j=1}^{n}\theta_{j}(\beta,f)x_{jk}\sum_{i=1}^{n}r_{ji}\frac{\delta_{i}}{s_{i}(\beta,f)}$
$=\sum_{i=1}^{n}\delta_{i}x_{ik}-\sum_{i=1}^{n}\theta_{i}(\beta,f)x_{ik}l_{i}(\beta,f)$
$=\sum_{i=1}^{n}(\delta_{i}-\theta_{i}(\beta,f)l_{i}(\beta,f))x_{ik},$
where $l_{i}(\beta,f)=\sum_{j=1}^{n}r_{ij}\frac{\delta_{j}}{s_{j}(\beta,f)}$. Continuing, we have:
$\frac{\partial\mathcal{L}(\beta,f)}{\partial\beta_{k}}$ $=\sum_{i=1}^{n}\frac{l_{i}(\beta,f)^{-1}\delta_{i}-\theta_{i}(\beta,f)}{\theta_{i}(\beta,f)l_{i}(\beta,f)^{-1}}\theta_{i}(\beta,f)x_{ik}$
$=\sum_{i=1}^{n}\frac{l_{i}(\beta,f)^{-1}\delta_{i}-\theta_{i}(\beta,f)}{\theta_{i}(\beta,f)l_{i}(\beta,f)^{-1}}\frac{\partial\theta_{i}(\beta,f)}{\partial\beta_{k}}.$
@@
$\frac{\partial\mathcal{L}(\beta,u)}{\partial f}=\sum_{i=1}^{n}\frac{l_{i}^{-1}\delta_{i}-\theta_{i}(\beta,f)}{v_{i}}\frac{\partial\theta_{i}(\beta,u)}{\partial f}.$
Thus, solving (12) is equivalent to finding $(\hat{\beta},\hat{f})$ that satisfies:
$\begin{cases}\sum_{i=1}^{n}\frac{l_{i}^{-1}\delta_{i}-\theta_{i}(\beta,f)}{v_{i}}\frac{\partial\theta_{i}(\beta,f)}{\partial\beta_{k}}=0,\quad\forall k=1,\dots,q,\\ \sum_{i=1}^{n}\frac{l_{i}^{-1}\delta_{i}-\theta_{i}(\beta,f)}{v_{i}}\frac{\partial\theta_{i}(\beta,u)}{\partial f}+\mu m(u,\hat{f})=0,\quad\forall u\in\mathcal{F}.\end{cases}$ (15)
@@
$\mathcal{J}_{\mu}(\beta,f)=\|V^{-1/2}(\delta/l-\theta)\|^{2}+\mu m(f,f),$
where $\delta/l=(\delta_{1}/l_{1},\dots,\delta_{n}/l_{n})^{\top}$ and $V=\text{diag}(v_{1},\dots,v_{n})$. Assuming such quantities are constant suggests an iterative computation scheme.
Let $\theta^{(k)}$ be an estimate of $\theta(\beta,f)$ after $k$ iterations. We consider a first-order approximation of $\theta$ near the current value $\theta^{(k)}=(\beta^{(k)},f^{(k)})$:
$\theta(\beta,f)\approx\underbrace{\exp(X\beta^{(k)}+f^{(k)})}_{\theta^{(k)}}+\frac{\partial\theta(\beta,f)}{\partial\beta}(\beta-\beta^{(k)})+\frac{\partial\theta(\beta,f)}{\partial f}(f-f^{(k)}).$
The partial derivatives of $\theta$ with respect to both parameters are computed as follows. For the derivative with respect to $\beta$:
$\frac{\partial\theta_{i}^{(k)}}{\partial\beta_{j}}=\partial\theta_{i}^{(k)}x_{ij},$
@@
$\frac{\partial\theta(\beta,f)}{\partial\beta}=\Theta^{(k)}X,$
where $\Theta^{(k)}=\text{diag}(\theta^{(k)})$.
@@
$\frac{\partial\theta_{i}(\beta^{(k)},f^{(k)}+\tau(f-f^{(k)}))}{\partial f}=\theta_{i}(\beta^{(k)},f^{(k)})(f(p_{i})-f^{(k)}(p_{i})).$
Hence, the first-order approximation of $\mathcal{J}_{\mu}(\beta,f)$ in the neighborhood of $(\beta^{(k)},f^{(k)})$ is:
$\tilde{\mathcal{J}}^{(k)}_{\lambda}(\beta,f)=\left\|V^{-1/2}\left[\delta/l-\left(\theta^{(k)}+\Theta^{(k)}X(\beta-\beta^{(k)})+\Theta^{(k)}(f-f^{(k)})\right)\right]\right\|^{2}+\mu m(f,f)$
$=\left\|V^{-1/2}\Theta^{(k)}\left((\Theta^{(k)})^{-1}(\delta/l-\theta^{(k)})+X\beta^{(k)}+f^{(k)}-X\beta-f\right)\right\|^{2}+\mu m(f,f).$
Defining $z^{(k)}=(\Theta^{(k)})^{-1}(\delta/l-\theta^{(k)})+X\beta^{(k)}+f^{(k)}$, and letting ${W}^{(k)}$ be the $n\times n$ diagonal matrix with $i$th entry equal to $v_{i}^{-1}(\beta^{(k)},f^{(k)})\theta_{i}^{2}(\beta^{(k)},f^{(k)})$, we can rewrite:
$\tilde{\mathcal{J}}^{(k)}_{\lambda}(\beta,f)=\left\|({W}^{(k)})^{1/2}(z^{(k)}-X\beta-f)\right\|^{2}+\mu m(f,f).$ (16)

### v2

@@ 7 Empirical Application
In this section, we apply the proposed method to analyze the London Fire Brigade (LFB) response time data. The dataset is available in the R package spatsurv (Taylor and Rowlingson,, 2017) and is further explored in Taylor, (2017). The original data can also be accessed via the London Datastore Fire and Authority, (2015). Our analysis focuses on data from 2009, specifically on fires occurring in residential dwellings, as these are associated with the highest number of fatalities. The response variable of interest is the time taken for the first fire engine to arrive at the scene of a dwelling fire. We excluded observations where the response time was missing, and the resulting dataset contains no censored data.
In Figure 3, we display the observed locations of these incidents, highlighting the spatial distribution of fires across London. The blue dots represent the locations of the observed fires, while the grey boundary outlines the geographical area under study. As can be seen from the image, the area excludes the River Thames, which corresponds to a natural boundary that could potentially influence the arrival time of the first fire engine. Given the irregular path of the river, it was necessary to smooth the boundary of the domain. After smoothing the domain and excluding points that fall into the river or outside the region of interest, the dataset results in 6657 observed locations, though less than 0.01% of the points were omitted. Some observations had identical locations, and while different strategies could be used, such as jittering the points, we chose to randomly omit the duplicated locations. This led to a dataset of 5949 unique locations, resulting in the removal of approximately 10% of the data points, which still leaves a sufficiently large dataset for analysis.
Figure 3: Observed fire locations in London. The blue points represent the geographical locations of fire incidents in residential dwellings across the city. The red border outlines the area under study, excluding the River Thames as a natural boundary.
Figure 4: Distribution of observed response times of the first fire engine to reach fire incidents. The color intensity represents the density of the response times across different locations.
Figure 5: Fire station intensity map. The blue intensity highlights the estimated proximity of fire stations to the observed fire locations, with red contours indicating areas of higher station density.
It is reasonable to expect that the response time to a fire service call is influenced by the proximity of the nearest fire station. However, in practice, the closest station may already be responding to another incident. As explored in Taylor and Rowlingson, (2017), we include a scaled fire station intensity measure as a covariate in the model, which accounts for the proximity of fire stations to fire locations. A mesh was constructed based on the observed locations using first-order basis functions. To determine the smooth parameter $\mu\in\{1,4,7,\dots,100\}$, we applied the CV-PLD method discussed in Section 4.5, utilizing a 10-fold cross-validation approach. The optimal $\mu$ was found to be 46. The estimation algorithm terminates if the Euclidean norm difference between consecutive estimates of the spatial field is less than 0.01, with convergence typically occurring within 5 iterations. The estimated regression coefficient for the fire station density covariate is $\beta=0.0667$, with a standard deviation of $1.41\times 10^{-5}$, indicating a statistically significant effect at the 1% level.
In Figure 6, we present the estimated spatial field. The figure highlights how the river acts as a natural boundary, constraining the fire engines’ ability to respond promptly to incidents on the opposite side of its banks. This observation suggests that geographical features significantly impact response efficiency and should be carefully considered when interpreting spatial variations in the data.
In Figure 7, we provide a cross-section of the spatial field, highlighted as a vertical line in Figure 6. This cross-section shows the estimated value of the spatial field at longitude 535000 as a function of latitude. We observe a distinct change point around latitude 180000, corresponding to the river’s location, where the hazard level immediately drops by approximately 80%.
It is arguable that the river effect might be artificially induced by omitting the river space from the domain and that the presence of bridges along the boundary effectively mitigated issues related to water. While the choice of domain should ideally be guided by a priori knowledge about the specific problem, it would nonetheless be valuable to conduct a sensitivity analysis regarding domain selection. Although this analysis is beyond the scope of this paper, we plan to address it in future studies.
Figure 6: Estimated spatial field of fire engine response. The red contour outlines areas where geographical boundaries, particularly the River Thames, influence the response efficiency of fire engines. The vertical dotted line corresponds to the section considered in Figure 7
Figure 7: Estimated spatial field at Latitude 535000, where it is possible to observe a change point at Latitude 180000 corresponding to the presence of the river.
@@ 8 Conclusions
This paper introduces a novel SPH model that integrates differential regularization and the finite element method to effectively address spatial dependencies in survival data observed over irregular domains. Our methodology expands on the traditional semiparametric proportional hazards model by incorporating a non-parametric spatial field component to account for spatial dependence, providing a flexible approach to capture unobserved geographical heterogeneity in hazard rates across a spatial domain.
The SPH model framework offers several advantages over existing methods. First, our approach employs differential regularization, which controls the smoothness of the estimated spatial field, allowing for adaptation to complex spatial boundaries. Secondly, by applying the finite element method on triangulated spatial meshes, we achieve computationally efficient and accurate estimates, particularly in irregularly shaped regions where spatial dependencies do not follow straightforward geometric patterns.
The proposed model demonstrated promising results in both simulation and empirical studies. Simulations confirmed that the model yields accurate estimates of both the covariate effects and the spatial component across various sample sizes and levels of censoring.
In future work, we aim to enhance the model’s adaptability by investigating alternative penalty structures, including anisotropic differential penalties, to more accurately capture directional dependencies in specific spatial datasets. Additionally, further research will focus on the asymptotic properties of the estimators to strengthen the model’s theoretical foundations. Sensitivity analysis concerning the choice of the boundary domain will also be pursued. Finally, we plan to extend the model to incorporate time-varying covariates within the spatial proportional hazards framework, enabling application to more dynamic spatial survival contexts, such as real-time risk prediction and event-driven spatial analysis.
@@ Appendix A Functional PIRLS Algorithm details
The maximizer $(\hat{\bm{\beta}},\hat{f})$ of (4) must satisfy the following system of first-order equations:
$\begin{cases}\frac{\partial\mathcal{L}(\hat{\bm{\beta}},\hat{f})}{\partial\beta_{k}}=0,\quad\forall k=1,\dots,q,\\ \lim_{\tau\to 0}\frac{1}{\tau}\left[\mathcal{L}(\hat{\bm{\beta}},\hat{f}+\tau u)-\mathcal{L}(\hat{\bm{\beta}},\hat{f})\right]-\mu m(u,\hat{f})=0,\quad\forall u\in\mathcal{F}.\end{cases}$ (12)
This system involves derivatives with respect to both parameters $\bm{\beta}$ and $f$, where the derivative with respect to $f$ is the Gâteaux derivative in the direction of $u\in\mathcal{F}$.
For each $k=1,\dots,q$, the term $\frac{\partial\mathcal{L}(\hat{\bm{\beta}},\hat{f})}{\partial\beta_{k}}$ is given by:
$\frac{\partial\mathcal{L}(\hat{\bm{\beta}},\hat{f})}{\partial\beta_{k}}=\sum_{i:\delta_{i}=1}\left(x_{ik}-\frac{\sum_{j:y_{j}\geq y_{i}}\exp\left(\bm{x}_{j}^{\top}\hat{\bm{\beta}}+\hat{f}(p_{j})\right)\bm{x}_{jk}}{\sum_{j:y_{j}\geq y_{i}}\exp\left(\bm{x}_{j}^{\top}\hat{\bm{\beta}}+\hat{f}(p_{j})\right)}\right).$ (13)
@@
$\mathcal{L}(\hat{\bm{\beta}},\hat{f}+\tau u)-\mathcal{L}(\hat{\bm{\beta}},\hat{f})$ $=\sum_{i:\delta_{i}=1}\left(tu(p_{i})-\log\frac{\sum_{j:y_{j}\geq y_{i}}\exp\left(\bm{x}_{j}^{\top}\hat{\bm{\beta}}+\hat{f}(p_{i})+tu(p_{i})\right)}{\sum_{j:y_{j}\geq y_{i}}\exp\left(\bm{x}_{j}^{\top}\hat{\bm{\beta}}+\hat{f}(p_{j})\right)}\right)$
$=\sum_{i:\delta_{i}=1}\left(tu(p_{i})-\log\frac{\sum_{j:y_{j}\geq y_{i}}\exp\left(\bm{x}_{j}^{\top}\hat{\bm{\beta}}+\hat{f}(p_{j})\right)(1+tu(p_{i})+o(t))}{\sum_{j:y_{j}\geq y_{i}}\exp\left(\bm{x}_{j}^{\top}\hat{\bm{\beta}}+\hat{f}(p_{j})\right)}\right)$
$=\sum_{i:\delta_{i}=1}\left(tu(p_{i})-\log\left\{1+t\frac{\sum_{j:y_{j}\geq y_{i}}\exp\left(\bm{x}_{j}^{\top}\hat{\bm{\beta}}+\hat{f}(p_{j})\right)u(p_{i})}{\sum_{j:y_{j}\geq y_{i}}\exp\left(\bm{x}_{j}^{\top}\hat{\bm{\beta}}+\hat{f}(p_{j})\right)}+o(t)\right\}\right)$
$=\sum_{i:\delta_{i}=1}\left(tu(p_{i})-t\frac{\sum_{j:y_{j}\geq y_{i}}\exp\left(\bm{x}_{j}^{\top}\hat{\bm{\beta}}+\hat{f}(p_{j})\right)u(p_{i})}{\sum_{j:y_{j}\geq y_{i}}\exp\left(\bm{x}_{j}^{\top}\hat{\bm{\beta}}+\hat{f}(p_{j})\right)}+o(t)\right).$
@@
$\lim_{\tau\to 0}\frac{1}{\tau}\left[\mathcal{L}(\hat{\bm{\beta}},\hat{f}+\tau u)-\mathcal{L}(\hat{\bm{\beta}},\hat{f})\right]$ $=\sum_{i:\delta_{i}=1}\left(u(p_{i})-\frac{\sum_{j:y_{j}\geq y_{i}}\exp\left(\bm{x}_{j}^{\top}\hat{\bm{\beta}}+\hat{f}(p_{j})\right)u(p_{i})}{\sum_{j:y_{j}\geq y_{i}}\exp\left(\bm{x}_{j}^{\top}\hat{\bm{\beta}}+\hat{f}(p_{j})\right)}\right).$ (14)
@@
$\theta_{i}(\bm{\beta},f)$ $=\exp(\bm{x}_{i}^{\top}\bm{\beta}+f(p_{i})),$
@@
$s_{i}(\bm{\beta},f)$ $=\sum_{j=1}^{n}r_{ji}\theta_{j}(\bm{\beta},f).$
@@
$\frac{\partial\mathcal{L}(\bm{\beta},f)}{\partial\beta_{k}}$ $=\sum_{i=1}^{n}\delta_{i}x_{ik}-\sum_{i=1}^{n}\frac{\delta_{i}}{s_{i}(\bm{\beta},f)}\sum_{j=1}^{n}r_{ji}\theta_{j}(\bm{\beta},f)\bm{x}_{jk}$
$=\sum_{i=1}^{n}\delta_{i}x_{ik}-\sum_{j=1}^{n}\theta_{j}(\bm{\beta},f)\bm{x}_{jk}\sum_{i=1}^{n}r_{ji}\frac{\delta_{i}}{s_{i}(\bm{\beta},f)}$
$=\sum_{i=1}^{n}\delta_{i}x_{ik}-\sum_{i=1}^{n}\theta_{i}(\bm{\beta},f)x_{ik}l_{i}(\bm{\beta},f)$
$=\sum_{i=1}^{n}(\delta_{i}-\theta_{i}(\bm{\beta},f)l_{i}(\bm{\beta},f))\bm{x}_{ik},$
where $l_{i}(\bm{\beta},f)=\sum_{j=1}^{n}r_{ij}\frac{\delta_{j}}{s_{j}(\bm{\beta},f)}$. Continuing, we have:
$\frac{\partial\mathcal{L}(\bm{\beta},f)}{\partial\beta_{k}}$ $=\sum_{i=1}^{n}\frac{l_{i}(\bm{\beta},f)^{-1}\delta_{i}-\theta_{i}(\bm{\beta},f)}{\theta_{i}(\bm{\beta},f)l_{i}(\beta,f)^{-1}}\theta_{i}(\bm{\beta},f)x_{ik}$
$=\sum_{i=1}^{n}\frac{l_{i}(\bm{\beta},f)^{-1}\delta_{i}-\theta_{i}(\bm{\beta},f)}{\theta_{i}(\bm{\beta},f)l_{i}(\bm{\beta},f)^{-1}}\frac{\partial\theta_{i}(\bm{\beta},f)}{\partial\beta_{k}}.$
@@
$\frac{\partial\mathcal{L}(\bm{\beta},u)}{\partial f}=\sum_{i=1}^{n}\frac{l_{i}^{-1}\delta_{i}-\theta_{i}(\bm{\beta},f)}{v_{i}}\frac{\partial\theta_{i}(\bm{\beta},u)}{\partial f}.$
Thus, solving (12) is equivalent to finding $(\hat{\bm{\beta}},\hat{f})$ that satisfies:
$\begin{cases}\sum_{i=1}^{n}\frac{l_{i}^{-1}\delta_{i}-\theta_{i}(\bm{\beta},f)}{v_{i}}\frac{\partial\theta_{i}(\bm{\beta},f)}{\partial\beta_{k}}=0,\quad\forall k=1,\dots,q,\\ \sum_{i=1}^{n}\frac{l_{i}^{-1}\delta_{i}-\theta_{i}(\bm{\beta},f)}{v_{i}}\frac{\partial\theta_{i}(\bm{\beta},u)}{\partial f}+\mu m(u,\hat{f})=0,\quad\forall u\in\mathcal{F}.\end{cases}$ (15)
@@
$\mathcal{J}_{\mu}(\bm{\beta},f)=\|V^{-1/2}(\bm{\delta}/\bm{l}-\theta)\|^{2}+\mu m(f,f),$
where $\bm{\delta}/\bm{l}=(\delta_{1}/l_{1},\dots,\delta_{n}/l_{n})^{\top}$, $\bm{v}=(v_{1},\dots,v_{n})^{\top}$, $V=\text{diag}(\bm{v})$ and $diag$ denotes a diagonal matrix with the $\bm{v}$ elements along its main diagonal, and zeros elsewhere. Assuming such quantities are constant suggests an iterative computation scheme.
Let $\bm{\theta}^{(k)}$ be an estimate of $\theta(\bm{\beta},f)$ after $k$ iterations. We consider a first-order approximation of $\theta$ near the current value $\theta^{(k)}=(\bm{\beta}^{(k)},f^{(k)})$:
$\theta(\bm{\beta},f)\approx\underbrace{\exp(X\bm{\beta}^{(k)}+f^{(k)})}_{\theta^{(k)}}+\frac{\partial\theta(\bm{\beta},f)}{\partial\beta}(\bm{\beta}-\bm{\beta}^{(k)})+\frac{\partial\theta(\bm{\beta},f)}{\partial f}(f-f^{(k)}).$
The partial derivatives of $\theta$ with respect to both parameters are computed as follows. For the derivative with respect to $\bm{\beta}$:
$\frac{\partial\theta_{i}^{(k)}}{\partial\beta_{j}}=\theta_{i}^{(k)}x_{ij},$
@@
$\frac{\partial\bm{\theta}(\bm{\beta},f)}{\partial\bm{\beta}}=\Theta^{(k)}X,$
where $\Theta^{(k)}=\text{diag}(\bm{\theta}^{(k)})$.
@@
$\frac{\partial\theta_{i}(\bm{\beta}^{(k)},f^{(k)}+\tau(f-f^{(k)}))}{\partial f}=\theta_{i}(\bm{\beta}^{(k)},f^{(k)})(f(p_{i})-f^{(k)}(p_{i})).$
Hence, the first-order approximation of $\mathcal{J}_{\mu}(\bm{\beta},f)$ in the neighborhood of $(\bm{\beta}^{(k)},f^{(k)})$ is:
$\tilde{\mathcal{J}}^{(k)}_{\lambda}(\beta,f)=\left\|V^{-1/2}\left[\bm{\delta}/\bm{l}-\left(\bm{\theta}^{(k)}+\Theta^{(k)}X(\bm{\beta}-\bm{\beta}^{(k)})+\Theta^{(k)}(f-f^{(k)})\right)\right]\right\|^{2}+\mu m(f,f)$
$=\left\|V^{-1/2}\Theta^{(k)}\left((\Theta^{(k)})^{-1}(\bm{\delta}/\bm{l}-\bm{\theta}^{(k)})+X\bm{\beta}^{(k)}+f^{(k)}-X\bm{\beta}-f\right)\right\|^{2}+\mu m(f,f).$
Defining $\bm{z}^{(k)}=(\Theta^{(k)})^{-1}(\bm{\delta}/\bm{l}-\theta^{(k)})+X\bm{\beta}^{(k)}+f^{(k)}$, and letting ${W}^{(k)}$ be the $n\times n$ diagonal matrix with $i$th entry equal to $v_{i}^{-1}(\bm{\beta}^{(k)},f^{(k)})\theta_{i}^{2}(\bm{\beta}^{(k)},f^{(k)})$, we can rewrite:
$\tilde{\mathcal{J}}^{(k)}_{\lambda}(\bm{\beta},f)=\left\|({W}^{(k)})^{1/2}(\bm{z}^{(k)}-X\bm{\beta}-f)\right\|^{2}+\mu m(f,f).$ (16)